3.3.12 Suppose you flip four fair coins. Let X be the number of heads showing, and
let Y be the number of tails showing. Compute Cov(X, Y) and Corr(X, Y).
1
Expert's answer
2020-12-17T18:46:15-0500

Solution

Y~Bin(4,0.5) and X~Bin(4,0.5)

Y=4-X

E(X)=np= 4(0.5)=2

E(Y)=np=4(0.5)=2

Var(X)=np(1-p)=1

Var(Y)=1

"Var(X) = E(x^2)-{E(x)}^2"

"E(x^2)=1+2^2=5"


"Cov(XY) =E(XY) - E(X) E(Y)""=E(X(4-X))-(2)(2)""=4E(X)-E(X^2)-4""=4(2)-5-4=-1"

Cov(XY) =-1


Corr(X, Y)


"Corr(XY) ={Cov(XY) \over \sigma_x \sigma_y}""={-1 \over (1*1)} =-1"

Corr(X, Y) =-1